how could we know the dimensions of complicated formulas from wave optics, magnetism,electricity etc?

Grade:12

1 Answers

pratham ashish
14 years ago
hi, solving for the dimension of a complicated formula has been always a tough task aomong the student community, and the situation get worse if u hav the time constraint but, DON'T WORRY v r here to give the proper technique for these type of probs, LOOK, ......... der some short tricks by which u can make yr lie easy such : if formula hav hc/lamda then consider the whole thing together & remember its dimension is that of energy (ML^2t^-2), similarly v*i hav power dimension, u have search these type of terms if not exactly present then make some manipulations by multiplying or dividing by some known thing, remember( mu*epsilon = c^2 dimension = M^2T^-2 ) for the time being learn to manipulate smwat, even if u face any prob then u can pist specific prob i wud elaborate the trick in more detail.